# Observational constraints on the Kerr and its several single-parameter modified spacetimes using quasi-periodic oscillation data

### v1

@@ I Introduction
[…]
@@
To date, numerous researches indicate that most celestial objects in the universe exhibit rotational characteristics a-1 ; a-2 ; a-3 . In GR, the unique stationary, axisymmetric, asymptotically flat vacuum solution to Einstein’s field equation is represented by the Kerr metric. This solution describes a non-charged rotating BH through two parameters, the mass $M$ and the spin parameter $a$ Kerr-1 ; Kerr-2 . However, the hypothesis that astrophysical BH candidates are characterized by Kerr spacetime still lacks sufficient evidence MG-aaa-1 . Although the predictions of GR regarding the existence of black holes have been confirmed, providing support for its validation in strong gravitational fields, the geometrical structure in strong gravitational fields for GR and the nonlinear behavior of modified gravity theories remain difficult to test and fully understand Kerr-3 . To address this, many modified forms of the Kerr BH have been proposed to expand our understanding of rotating black holes. Examples include regular BH models Bardeen-1 ; ABG-1 ; Hayward-1 , BH solutions with a charge KN-1 ; KTN-1 ; BK-1 , BH solutions obtained under different modified gravity theories MK-1 ; KS-1 , and BH solutions considering dark matter effects PFDM-1 ; PFDM-2 etc. These modified forms of the Kerr spacetime provide new perspectives for the study of BH geometry and physical properties, opening new research directions in astrophysics. In recent years, astronomical observations, such as quasi-periodic oscillations (QPOs) detected from microquasars QPO-1 ; QPO-2 ; QPO-3 ; QPO-4 ; QPO-5 , the shadow of the Sgr A* BH observed by telescopes shadow-1 ; shadow-2 , and flare hotspot data observed by gravitational instruments hot-1 , have become effective ways to test the different gravitational theories.
@@
QPOs are one of the most important tools for testing gravitational theories due to their high precision characteristics NS-8 ; NS-9 ; QPO-1 ; QPO-2 ; QPO-3 . QPOs refer to the peaks observed in the electromagnetic spectrum, from the radio to X-ray bands, originating from compact objects QPO-6 . Based on the observed frequencies, QPOs are classified into low-frequency (LF) QPOs and high-frequency (HF) QPOs. By analyzing the spectral characteristics of QPOs, scientists can extract physical information about the central object QPO-6 ; QPO-8 ; QPO-9 ; QPO-10 . Although the generation mechanism of QPOs is not yet fully understood, they are generally believed to be related to relativistic precession and resonance phenomena associated with GR effects QPO-11 ; QPO-12 ; QPO-13 . To investigate these phenomena, various QPO models, including those based on relativistic precession and resonance, have been proposed QPO-model-1 ; QPO-model-2 ; QPO-model-3 ; QPO-model-4 ; QPO-model-5 ; QPO-model-6 ; QPO-model-7 ; QPO-model-8 . The construction of these models typically involves the motion of particles around the central object and considers the linear combinations of radial, latitudinal, and orbital frequencies. This paper primarily investigates the dynamical behavior of particles in the Kerr spacetime and nine single-parameter modified versions of the Kerr geometry (Bardeen, Ayon-Beato and García (ABG), Hayward, Kerr-Newman (KN), Kerr-Taub-NUT (KTN), Braneworld Kerr (BK), Kerr-modified theory of gravity (kerr-MOG), Kerr-Sen, and Perfect Fluid Dark Matter (PFDM) black holes), testing these spacetimes using QPO observational data from three microquasars (GRO J1655-40, H1743-322, XTE J1859+226).
@@ II Models of stationary axisymmetric spacetime
For a stationary axisymmetric spacetime with general properties, the expression in Boyer-Lindquist (BL) coordinates is written by BL :
$ds^{2}=g_{tt}dt^{2}+2g_{t\phi}d\phi dt+g_{\phi\phi}d\phi^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}.$ (1)
The non-zero metric components of the Kerr spacetime in BL coordinates are given by Kerr-4 ; Kerr-5 ,
@@
where $M$ is the mass of the BH, and $a$ is the spin parameter. Next, we will consider nine stationary, axisymmetric, asymptotically flat spacetime solutions. These solutions are single-parameter modifications of the Kerr spacetime, and when the modification parameter for each model is set to 0, they reduce to the Kerr geometry. These spacetimes encompass not only solutions within the framework of GR (such as rotating regular BH and rotating BH with a charge) but also BH solutions obtained by modified gravitational theories and those derived considering dark matter effects.
@@ II.1 Rotating regular BHs in GR
The regular BH was first introduced as a phenomenological model Bardeen-1 . In recent studies, spherically symmetric regular BHs have been found to serve as nonsingular exact solutions to the GR field equations coupled to nonlinear electrodynamics regular-s-1 ; regular-s-2 . The rotating regular BH, on the other hand, are derived from the spherically symmetric regular BH solution through the non-complexification process of the Newman-Janis algorithm NJ-1 . Next, we will introduce three classical rotating regular BH solutions within the framework of GR.
@@
Ayón-Beato and Garcia (ABG) ABG-1 ; ABG-2 ; ABG-3 also obtained a BH solution free of singularity, referred as the ABG BH. Unlike the Bardeen BH, the ABG BH is a rigorous mathematical solution within the framework of GR coupled with nonlinear electrodynamics. In BL coordinates, the metric components of the axisymmetric form of the ABG BH are given by ABG-1 ; NJ-1
@@
where ${}F=1-\frac{2Mr^{2}}{\left(r^{2}+Q_{1}{}^{2}\right)^{3/2}}+\frac{Q_{1}{}^{2}r^{2}}{\left(r^{2}+Q_{1}{}^{2}\right)^{2}}$, and $Q_{1}$ represents the charge. When $Q_{1}=0$, (4) is equivalent to classical Kerr solution.
@@ II.3 Rotating BHs in modified theories of gravity
The Kerr-MOG BH, derived by Moffat based on the MOG theory, is a rotating black hole solution MK-1 . MOG gravity, also called as scalar-tensor-vector gravity (STVG), was proposed by Moffat with the aim of explaining gravitational phenomena in galaxies and the universe without the need for dark matter assumptions MK-1 . This MOG is based on an action, which consists of the usual Einstein-Hilbert term associated with the metric $g_{\mu\nu}$, a massive vector field $\varphi_{\mu}$, and three scalar fields that represent the running values of the gravitational constant $G$, the vector field mass $\mu$, and its coupling strength $\omega$, respectively MK-2 . The geometry of the Kerr-MOG BH can be expressed by the line element in the standard BL coordinates MK-2 ; MK-3
$\begin{gathered}g_{tt}=-\left(\frac{\Delta_{7}-a^{2}\sin^{2}\theta}{\Sigma}\right),\quad g_{t\phi}=\frac{a\sin^{2}\theta}{\Sigma}\left[\Delta_{7}-\left(r^{2}+a^{2}\right)\right],\quad g_{\theta\theta}=\Sigma,\\ g_{\phi\phi}=\frac{\sin^{2}\theta}{\Sigma}\left[\left(r^{2}+a^{2}\right)^{2}-\Delta_{7}a^{2}\sin^{2}\theta\right],\quad g_{rr}=\frac{\Sigma}{\Delta_{7}},\\ \Delta_{7}=r^{2}-2G_{\mathrm{N}}Mr+a^{2}+\alpha GG_{\mathrm{N}}M^{2},\quad\Sigma=r^{2}+a^{2}\cos^{2}\theta,\end{gathered}$ (9)
where $G_{\mathrm{N}}=G(1+\alpha)$ is the enhanced gravitational constant. In the subsequent calculations, we choose the gravitational constant to be $G=1$. $\alpha$ is the modification parameter, and when $\alpha=0$ the Kerr-MOG solution reduces to Kerr geometry.
@@ III The dynamical effects of particles in some stationary axisymmetric spacetimes
In the spacetime characterized by equation (1), the geodesic motion of a particle near the central body is governed by $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=\epsilon$. When $\epsilon=0$, it corresponds to null geodesics, and when $\epsilon=-1$, it corresponds to timelike geodesics. By using the four-dimensional normalization equation, for a massive particle moving along a timelike geodesic in the equatorial plane, the following can be acquired
@@
The symbols +/+/-+ / - represent the angular velocity of particles undergoing prograde and retrograde rotations, respectively. To determine the angular frequencies of the radial and latitudinal epicyclic motion of the particle on a stable circular orbit, we consider the geodesic equation w-1
@@
Figure 1: With a spin parameter $a=0.5$, in the Kerr and its nine types of single-parameter modified spacetimes, the frequencies of the test particle, $\omega_{r}$ (dotted line), $\omega_{\theta}$ (dashed line), and $\omega_{\phi}$ (solid line), are plotted as functions of the radial coordinate $r$.
In Figure 1, we present the variation of the test particle’s frequencies $\omega_{r}$ (dotted line), $\omega_{\theta}$ (dashed line), and $\omega_{\phi}$ (solid line) with respect to the radial coordinate $r$ in the Kerr spacetime, as well as in the nine types of single-parameter modified spacetimes of Kerr. In the plot, we take spin parameter $a=0.5$, and only the range $r>r_{\text{ISCO }}$ is considered. Additionally, to visually illustrate the impact of the modification parameters on the test particle’s frequencies, we set the modification parameter for all nine types of modified Kerr spacetimes to 0.5. From the figure, it is observed that as the particle approaches the BH, its radial frequency distribution in the non-Kerr spacetimes significantly deviates from that in the Kerr spacetime. However, as the particle moves away from the black hole, the radial frequency distributions across the different spacetimes gradually converge. To date, the particle’s epicyclic motion and orbital motion around the central body are widely regarded as the primary physical sources of the observed QPO phenomenon. Researchers focus on using the angular frequencies of the particle’s epicyclic motion and orbital motion to construct corresponding theoretical models, and by combining these with the observed QPO phenomena, conduct in-depth studies of the astrophysical QPO behavior. To ensure consistency between the physical quantities in the theoretical model and the observational data in terms of dimensional analysis, we introduce the following definition
@@ IV Data fitting and observational constraints on models of stationary axisymmetric spacetime
In the study of QPO phenomena, the construction of related models typically involves linear combinations of the radial, latitudinal epicyclic motion frequencies, and orbital frequencies of a particle’s around the central body. Overall, QPO models can be broadly classified into four categories: the tidal precession model QPO-model-9 , the disk oscillation model QPO-model-10 ; QPO-model-11 ; QPO-model-12 , the relativistic precession model QPO-model-13 ; QPO-model-14 , and the resonance model QPO-model-1 . Considering the QPO data observed in the three aforementioned microquasars, to date, only the RP model has been able to simultaneously explain the effects of both HFQPO and LFQPO LFQPO-1 ; LFQPO-2 . Therefore, this paper will employ the RP model to analyze the QPO observational data presented in Table 1.
@@
According to the calculations listed in Table 2, we can find some interesting results on the constraints on modification parameters: (i) The constraints on modification parameters $\Delta^{*}$ for all axisymmetric spacetimes investigated are strict by using the QPO data, i.e. at the $68\%$ CL they have the relative narrow confidence intervals; (ii) At the $68\%$ CL, only the constraint on modification parameter $Q_{2}^{*}$ in the KN spacetime admits both negative and positive modifications (where zero recovers the Kerr solution), while the confidence ranges of modification parameters for other geometries require all strictly positive; (iii) According to the constraint results derived from QPO data, the small values of modification parameters are suggested in the KN and PFDM geometries ($Q_{2}^{*}=-0.004\pm 0.078$ and $k^{*}=$ $0.019_{-0.003}^{+0.024}$), while the larger value of modification parameter is favored in Kerr-MOG geometry ($\alpha^{*}=0.795\pm 0.011$); (iv) The limit results on $\Delta^{*}$ in all modified version of Kerr (except KN model) do not include the value $\Delta^{*}=0$ (corresponding to Kerr) at the $68\%$ CL, which indicates that the investigation into modifications of the Kerr spacetime is valuable; (v) Although the modification parameters in the ABG, KN, and Kerr-Sen spacetimes are all related to electric charge, they influence geometry in different ways, which may lead to different constraint results on the parameters. Concretely, the constraints on $Q_{1}^{*}=0.360_{-0.011}^{+0.012}$ and $Q_{3}^{*}=0.361_{-0.024}^{+0.028}$ are similar, while they differ significantly from that of $Q_{2}^{*}=-0.004\pm 0.078$; (vi) The modification parameters in Bardeen and KTN geometries are interpreted as the effect originating from the magnetic monopole. It is shown in Table 2 that $b^{*}=0.229_{-0.034}^{+0.045}$ and $n^{*}=0.257_{-0.025}^{+0.054}$ exhibit closely matching values.
@@
Additionally, the constraints on some modification parameters in the single-parameter modified spacetimes of Kerr have been discussed in several other studies. For instance, in the study by Bardeen-2 , the authors used QPO observational data from microquasars (GRO J1655-40, XTE J1550-564, GRS 1915+105, H1743-322) and the $\mathrm{Sgr~{}A}^{*}$ data to constrain the Bardeen spacetime. The best fit, which minimizes $\chi^{2}$, was found for $b^{*}\cong 0.25$ under the RP model, which is consistent with our constraint: $b^{*}=0.229_{-0.034}^{+0.045}$; In Ref. modified-Kerr-1 , the authors combined the LIGO-Virgo observational data of the GW150914 event with the KN spacetime and derived a constraint on the charge $Q_{2}^{*}$ within the $90\%$ CL as $Q_{2}^{*}<0.33$. This result is consistent with our constraint on $Q_{2}^{*}$ in the KN spacetime, which is $-0.004\pm 0.078$; In the study by modified-Kerr-2 , the authors combined the KTN spacetime with observational data, including the radiation efficiency and jet power of the microquasar GR0 J1655-40, to determine the NUT parameter $n^{*}$ range as $0.80<n^{*}<$ 0.8. Our result for the NUT parameter $n^{*}=0.257_{-0.025}^{+0.054}$ falls within this range, and has a more strict constraint; In modified-Kerr-3 , the authors employed the parameter resonance model to combine the BK spacetime with HFQPO observational data from the microquasar GRS 1915+105, resulting in a constraint on the parameter $q^{*}$ within the range $[-3.258,0.287]$. This range has a more large error than our result and includes the constraint on $q^{*}$ derived in this work, which is $0.152\pm 0.019$; In modified-Kerr-4 , the authors considered the spin parameter $a^{*}$ and the geometric structure of the Kerr-MOG BH, and received a constraint on the deformation parameter $0\leq\alpha^{*}\leq 5.250$ in theory when the spin parameter $a^{*}\approx 0.4$. This range encompasses the constraint on the deformation parameter $\alpha^{*}$ obtained in our study, which is $0.795\pm 0.011$; In modified-Kerr-5 , the authors incorporated the PFDM spacetime with the Event Horizon Telescope (EHT) observations of M87*, determining a constraint on the parameter $k^{*}$ within the $1\sigma$ confidence region, which spans the range $[0,0.0792]$. This result is in agreement with the constraint we derived for the PFDM parameter $k^{*}$, which is $0.019_{-0.003}^{+0.002}$.
@@
Different from the consistency between the mass ranges constrained by using the QPO method and those derived from optical/near-infrared luminosity measurements, the spin values for the three microquasars derived using the QPO method across the ten models exhibit substantial discrepancies compared to the spin values obtained through the iron line/continuum-fitting method in Kerr spacetime. For instance, for the microquasar GRO J1655-40, the spin value $a^{*}$ derived from the continuumfitting method is $0.7\pm 0.1$ cf . In contrast, in all ten models examined in our study, constraints on the spin value for this microquasar, derived using the QPO method within the $68\%$ CL is below 0.5. This discrepancy has also been found in some other literature, e.g. with possible explanations proposed in GRO-1 ; GRO-2 ; GRO-3 ; GRO-4 . Nonetheless, this difference remains an area worthy of further investigation and exploration.
Finally, the constraint results for various parameters in ten different rotationally symmetric spacetimes, as presented in this study, indicate that the QPO method could provide stricter constraints on spin values, mass, and modification parameters compared to methods like optical/near-infrared luminosity measurements. Furthermore, differences in the selected model can significantly influence the final results, making the QPO method a powerful tool for testing gravitational theories. And it is meaningful to assess the status of various spacetimes, when fitting the observed data to constrain the parameters in physics.

### v2

@@ I Introduction
The formulation of General Relativity (GR) has profoundly impacted on the development of physics. In weak gravitational fields, the predictions of GR in the solar system and beyond are consistent with most astronomical observations GR-1 , such as light deflection GR-2 , gravitational redshift and the correction to the perihelion precession of mercury GR-3 , etc.. Recently, observations such as gravitational wave signals from binary black hole (BH) mergers GR-4 and the shadow image of the supermassive black hole $\mathrm{M87^{*}}$ GR-5 ; GR-6 ; GR-7 ; Letter-11 have further strongly supported GR’s predictions regarding the existence of BHs.
@@
To date, numerous researches indicate that most celestial objects in the universe exhibit rotational characteristics a-1 ; a-2 ; a-3 . In GR, the unique stationary, axisymmetric, asymptotically flat vacuum solution to Einstein’s field equation is represented by the Kerr metric. This solution describes a non-charged rotating BH through two parameters, the mass $M$ and the spin parameter $a$ Kerr-1 ; Kerr-2 . However, the hypothesis that astrophysical BH candidates are characterized by Kerr spacetime still lacks sufficient evidence MG-aaa-1 . Although the predictions of GR regarding the existence of black holes have been confirmed, providing support for its validation in strong gravitational fields, the geometrical structure in strong gravitational fields for GR and the nonlinear behavior of modified gravity theories remain difficult to test and fully understand Kerr-3 . To address this, many modified forms of the Kerr BH have been proposed to expand our understanding of rotating black holes. Examples include regular BH models Bardeen-1 ; ABG-1 ; Hayward-1 , BH solutions with a charge KN-1 ; KTN-1 ; BK-1 , BH solutions obtained under different modified gravity theories MK-1 ; KS-1 , and BH solutions considering dark matter effects PFDM-1 ; PFDM-2 etc. These modified forms of the Kerr spacetime provide new perspectives for the study of BH geometry and physical properties, opening new research directions in astrophysics. In recent years, astronomical observations, such as quasi-periodic oscillations (QPOs) detected from microquasars QPO-1 ; QPO-2 ; QPO-3 ; QPO-4 ; QPO-5 , the shadow of the Sgr A* BH observed by telescopes shadow-1 ; shadow-2 ; Letter-1 ; Letter-4 ; Letter-5 ; Letter-5.5 , gravitational waves Letter-2 ; Letter-3 ; Letter-6 and flare hotspot data observed by gravitational instruments hot-1 , have become effective ways to test the different gravitational theories.
@@
QPOs are one of the most important tools for testing gravitational theories due to their high precision characteristics NS-8 ; NS-9 ; QPO-1 ; QPO-2 ; QPO-3 . QPOs phenomenon can be observed in the X-ray flux from black hole and neutron star X-ray binary systems, and detected as narrow peaks in the power density spectrum LFQPO-1 ; QPO-6-2 . Based on the observed frequencies, QPOs are classified into low-frequency (LF) QPOs and high-frequency (HF) QPOs. By analyzing the spectral characteristics of QPOs, scientists can extract physical information about the central object QPO-6 ; QPO-8 ; QPO-9 ; QPO-10 . Although the generation mechanism of QPOs is not yet fully understood, they are generally believed to be related to relativistic precession and resonance phenomena associated with GR effects QPO-11 ; QPO-12 ; QPO-13 . To investigate these phenomena, various QPO models, including those based on relativistic precession and resonance, have been proposed QPO-model-1 ; QPO-model-2 ; QPO-model-3 ; QPO-model-4 ; QPO-model-5 ; QPO-model-6 ; QPO-model-7 ; QPO-model-8 . The construction of these models typically involves the motion of particles around the central object and considers the linear combinations of radial, latitudinal, and orbital frequencies. This paper primarily investigates the dynamical behavior of particles in the Kerr spacetime and nine single-parameter modified versions of the Kerr geometry (Bardeen, Ayon-Beato and García (ABG), Hayward, Kerr-Newman (KN), Kerr-Taub-NUT (KTN), Braneworld Kerr (BK), Kerr-modified theory of gravity (kerr-MOG), Kerr-Sen, and Perfect Fluid Dark Matter (PFDM) black holes), testing these spacetimes using QPO observational data from three microquasars (GRO J1655-40, H1743-322, XTE J1859+226).
@@ II Models of stationary axisymmetric spacetime
Let us consider the line element of a circular, stationary, and axisymmetric spacetime written in canonical form BL :
$ds^{2}=g_{tt}dt^{2}+2g_{t\phi}d\phi dt+g_{\phi\phi}d\phi^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2},$ (1)
where there is only one non-vanishing off-diagonal metric coefficient. The non-zero metric components of the Kerr spacetime in Boyer-Lindquist (BL) coordinates are given by Kerr-4 ; Kerr-5 ,
@@
where $M$ is the mass of the BH, and $a$ is the spin parameter. Next, we will consider nine stationary, axisymmetric, asymptotically flat spacetime solutions. These solutions are single-parameter modifications of the Kerr spacetime, and when the modification parameter for each model is set to 0, they reduce to the Kerr geometry. These spacetimes encompass not only solutions within the framework of GR (such as rotating regular BH and rotating BH with a charge) but also BH solutions obtained by modified gravitational theories and those derived considering dark matter effects. In this section we use geometrized units.
@@ II.1 Rotating regular BHs in GR
The regular BH was first introduced as a phenomenological model Bardeen-1 . In recent studies, spherically symmetric regular BHs have been found to serve as nonsingular exact solutions to the GR field equations coupled to nonlinear electrodynamics regular-s-1 ; regular-s-2 . The rotating regular BH, on the other hand, are derived from the spherically symmetric regular BH solution through the non-complexification process NJ-1 . Next, we will introduce three classical rotating regular BH solutions within the framework of GR.
@@
Ayón-Beato and Garcia (ABG) ABG-1 ; ABG-2 ; ABG-3 also obtained a BH solution free of singularity, referred as the ABG BH. Unlike the Bardeen BH, the ABG BH is a rigorous mathematical solution within the framework of GR coupled with nonlinear electrodynamics. The metric components of the axisymmetric form of the ABG BH are given by ABG-1 ; NJ-1
@@
where $F=1-\frac{2Mr^{2}}{\left(r^{2}+Q_{1}{}^{2}\right)^{3/2}}+\frac{Q_{1}{}^{2}r^{2}}{\left(r^{2}+Q_{1}{}^{2}\right)^{2}}$ and $Q_{1}$ represents the charge. When $Q_{1}=0$, (4) is equivalent to classical Kerr solution.
@@ II.3 Rotating BHs in modified theories of gravity
The Kerr-MOG BH, derived by Moffat based on the MOG theory, is a rotating black hole solution MK-1 . MOG gravity, also called as scalar-tensor-vector gravity (STVG), was proposed by Moffat with the aim of explaining gravitational phenomena in galaxies and the universe without the need for dark matter assumptions MK-1 . This MOG is based on an action, which consists of the usual Einstein-Hilbert term associated with the metric $g_{\mu\nu}$, a massive vector field $\varphi_{\mu}$, and three scalar fields that represent the running values of the gravitational constant $G_{\mathrm{K}}$, the vector field mass $\mu$, and its coupling strength $\omega$, respectively MK-2 . The geometry of the Kerr-MOG BH can be expressed by the line element in the standard BL coordinates MK-2 ; MK-3
$\begin{gathered}g_{tt}=-\left(\frac{\Delta_{7}-a^{2}\sin^{2}\theta}{\Sigma}\right),\quad g_{t\phi}=\frac{a\sin^{2}\theta}{\Sigma}\left[\Delta_{7}-\left(r^{2}+a^{2}\right)\right],\quad g_{\theta\theta}=\Sigma,\\ g_{\phi\phi}=\frac{\sin^{2}\theta}{\Sigma}\left[\left(r^{2}+a^{2}\right)^{2}-\Delta_{7}a^{2}\sin^{2}\theta\right],\quad g_{rr}=\frac{\Sigma}{\Delta_{7}},\\ \Delta_{7}=r^{2}-2G_{\mathrm{K}}Mr+a^{2}+\alpha GG_{\mathrm{K}}M^{2},\quad\Sigma=r^{2}+a^{2}\cos^{2}\theta,\end{gathered}$ (9)
where $G_{\mathrm{K}}=G(1+\alpha)$ is the enhanced gravitational constant, where $G=1$ is Newtonian gravitational constant. $\alpha$ is the modification parameter, and when $\alpha=0$ the Kerr-MOG solution reduces to Kerr geometry.
@@ III The dynamical effects of particles in some stationary axisymmetric spacetimes
In the spacetime characterized by equation (1), the geodesic motion of a particle near the central body is governed by $g_{\mu\nu}\dot{x}^{\mu}\dot{x}^{\nu}=\epsilon$. When $\epsilon=0$, it corresponds to null geodesics, and when $\epsilon=-1$, it corresponds to timelike geodesics. For a massive particle moving along a timelike geodesic in the equatorial plane, the following relation can be acquired
@@
The symbols +/+/- represent the angular velocity of particles undergoing prograde and retrograde rotations, respectively. To determine the angular frequencies of the radial and latitudinal epicyclic motion of the particle on a stable circular orbit, we consider the geodesic equation w-1
@@
Figure 1: With a spin parameter $a=0.5$, in the Kerr and its nine types of single-parameter modified spacetimes (set each modification parameter to 0.5), the frequencies of the test particle, $\omega_{r}$ (dotted line), $\omega_{\theta}$ (dashed line), and $\omega_{\phi}$ (solid line), are plotted as functions of the radial coordinate $r$.
In Figure 1, we present the variation of the test particle’s frequencies $\omega_{r}$ (dotted line), $\omega_{\theta}$ (dashed line), and $\omega_{\phi}$ (solid line) with respect to the radial coordinate $r$ in the Kerr spacetime, as well as in the nine types of single-parameter modified spacetimes of Kerr. In the plot, we take spin parameter $a=0.5$, and only the range $r>r_{\text{ISCO }}$ is considered. Additionally, to visually illustrate the impact of the modification parameters on the test particle’s frequencies, we set the modification parameter for all nine types of modified Kerr spacetimes to 0.5. From the figure, it is observed that as the particle approaches the BH, its radial frequency distribution in the non-Kerr spacetimes significantly deviates from that in the Kerr spacetime. However, as the particle moves away from the black hole, the radial frequency distributions across the different spacetimes gradually converge. One knows that the particle’s epicyclic and orbital motions around the central body are theoretically regarded as one possible origin of the observed QPO phenomenon. Researchers focus on using the angular frequencies of the particle’s epicyclic motion and orbital motion to construct corresponding theoretical models, and by combining these with the observed QPO phenomena, conduct in-depth studies of the astrophysical QPO behavior. To facilitate the subsequent use of QPO observational data in constraining various models, we introduce
@@ IV Data fitting and observational constraints on models of stationary axisymmetric spacetime
In the study of QPO phenomena, the construction of related models typically involves linear combinations of the radial, latitudinal epicyclic motion frequencies, and orbital frequencies of a particle around the central body. Overall, QPO models can be broadly classified into four categories: the tidal precession model QPO-model-9 , the disk oscillation model QPO-model-10 ; QPO-model-11 ; QPO-model-12 , the relativistic precession model QPO-model-13 ; QPO-model-14 , and the resonance model QPO-model-1 . Considering the QPO data observed in the three aforementioned microquasars, to date, only the RP model has been able to simultaneously explain the effects of both HFQPO and LFQPO LFQPO-1 ; LFQPO-2 . Therefore, this paper will employ the RP model to analyze the QPO observational data presented in Table 1.
@@
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@
According to the calculations listed in Table 2, we can find some interesting results on the constraints on modification parameters: (i) The constraints on modification parameters $\Delta^{*}$ for all axisymmetric spacetimes investigated are strict by using the QPO data, i.e. at the $68\%$ CL they have the relative narrow confidence intervals; (ii) At the $68\%$ CL, only the constraint on modification parameter $Q_{2}^{*}$ in the KN spacetime admits both negative and positive modifications (where zero recovers the Kerr solution), while the confidence ranges of modification parameters for other geometries require all strictly positive; (iii) According to the constraint results derived from QPO data, the small values of modification parameters are suggested in the KN and PFDM geometries ($Q_{2}^{*}=-0.004\pm 0.078$ and $k^{*}=$ $0.019_{-0.003}^{+0.024}$). The nonzero constrained value of $k^{*}$ may provide a signature of dark matter, although its relatively small magnitude might suggest a weak effect; (iv) The larger value of modification parameter is favored in Kerr-MOG geometry ($\alpha^{*}=0.795\pm 0.011$), which indicates that the gravitational field of the Kerr-MOG black hole is stronger than that of the traditional Kerr solution, potentially exhibiting more significant observational features; (v) The limit results on $\Delta^{*}$ in all modified version of Kerr (except KN model) do not include the value $\Delta^{*}=0$ (corresponding to Kerr) at the $68\%$ CL, which indicates that the investigation into modifications of the Kerr spacetime is valuable; (vi) Although the modification parameters in the ABG, KN, and Kerr-Sen spacetimes are all related to electric charge, they influence geometry in different ways, which may lead to different constraint results on the parameters. Concretely, the constraints on $Q_{1}^{*}=0.360_{-0.011}^{+0.012}$ and $Q_{3}^{*}=0.361_{-0.024}^{+0.028}$ are similar, while they differ significantly from that of $Q_{2}^{*}=-0.004\pm 0.078$; (vii) The modification parameters in Bardeen and KTN geometries are interpreted as the effect originating from the magnetic monopole. It is shown in Table 2 that $b^{*}=0.229_{-0.034}^{+0.045}$ and $n^{*}=0.257_{-0.025}^{+0.054}$ exhibit closely matching values.
@@
Additionally, the constraints on some modification parameters in the single-parameter modified spacetimes of Kerr have been discussed in several other studies. For instance, in the study by Bardeen-2 , the authors used QPO observational data from microquasars (GRO J1655-40, XTE J1550-564, GRS 1915+105, H1743-322) and the $\mathrm{Sgr~A}^{*}$ data to constrain the Bardeen spacetime. The best fit, which minimizes $\chi^{2}$, was found for $b^{*}\cong 0.25$ under the RP model, which is consistent with our constraint: $b^{*}=0.229_{-0.034}^{+0.045}$; In Ref. modified-Kerr-1 , the authors combined the LIGO-Virgo observational data of the GW150914 event with the KN spacetime and derived a constraint on the charge $Q_{2}^{*}$ within the $90\%$ CL as $Q_{2}^{*}<0.33$. This result is consistent with our constraint on $Q_{2}^{*}$ in the KN spacetime, which is $0.004\pm 0.078$; In the study by modified-Kerr-2 , the authors combined the KTN spacetime with observational data, including the radiation efficiency and jet power of the microquasar GR0 J1655-40, to determine the NUT parameter $n^{*}$ range as $0<n^{*}<0.8$. Our result for the NUT parameter $n^{*}=0.257_{-0.025}^{+0.054}$ falls within this range, and has a more strict constraint; In modified-Kerr-3 , the authors employed the parameter resonance model to combine the BK spacetime with HFQPO observational data from the microquasar GRS 1915+105, resulting in a constraint on the parameter $q^{*}$ within the range [3.258,0.287][-3.258,0.287]. This range has a more large error than our result and includes the constraint on $q^{*}$ derived in this work, which is $0.152\pm 0.019$; In modified-Kerr-4 , the authors considered the spin parameter $a^{*}$ and the geometric structure of the Kerr-MOG BH, and received a constraint on the deformation parameter $0\leq\alpha^{*}\leq 5.250$ in theory when the spin parameter $a^{*}\approx 0.4$. This range encompasses the constraint on the deformation parameter $\alpha^{*}$ obtained in our study, which is $0.795\pm 0.011$; In modified-Kerr-5 , the authors incorporated the PFDM spacetime with the Event Horizon Telescope (EHT) observations of M87*, determining a constraint on the parameter $k^{*}$ within the $1\sigma$ confidence region, which spans the range $[0,0.0792]$. This result is in agreement with the constraint we derived for the PFDM parameter $k^{*}$, which is $0.019_{-0.003}^{+0.002}$.
@@
Different from the consistency between the mass ranges constrained by using the QPO method and those derived from optical/near-infrared luminosity measurements, the spin values for the three microquasars derived using the QPO method across the ten models exhibit substantial discrepancies compared to the spin values obtained through the iron line/continuum-fitting method in Kerr spacetime. For instance, for the microquasar GRO J1655-40, the spin value $a^{*}$ derived from the continuumfitting method is $0.7\pm 0.1$ cf . In contrast, in all ten models examined in our study, constraints on the spin value for this microquasar, derived using the QPO method within the $68\%$ CL is below 0.5. This discrepancy has also been found in some other literature, e.g. with possible explanations proposed in GRO-1 ; GRO-2 ; GRO-3 ; GRO-4 . As pointed out in reference GRO-1 , if the RP model is reliable for characterizing the quasi-periodic oscillation behavior of celestial bodies, the inconsistency may stem from the overestimation of the spin parameter by the continuum-fitting method, and may even suggest a violation of the Kerr black hole paradigm. Also, this inconsistency could be attributed to the limitation of the QPO model used, as well as the misunderstanding or insufficient understanding of the physical processes in accretion flows. In conclusion, this difference remains an area worthy of further investigation and exploration.
Finally, the constraint results for various parameters in ten different rotationally symmetric spacetimes, as presented in this study, indicate that the QPO method could provide stricter constraints on spin values, mass, and modification parameters compared to methods like optical/near-infrared luminosity measurements. Furthermore, differences in the selected model can significantly influence the final results, making the QPO method a powerful tool for testing gravitational theories. And it is meaningful to assess the status of various spacetimes by combing the observed data with the theoretical models.